OpenStudy (anonymous):

Which sum or difference identity would you use to verify that cos (180° - q) = -cos q?

OpenStudy (butterflydreamer):

You want to use the cos difference rule: Cos ( A - B) = CosA* CosB + SinA* SinB

OpenStudy (anonymous):

what will (a-b) be

OpenStudy (butterflydreamer):

So in this case your A = 180 and B = q So simply substitute that into the above equation i wrote for you :) Cos ( 180 - q ) = Cos180 * Cos q + Sin180 * Sin q [note: Cos 180 = -1 and Sin 180 = 0 ] This would mean : Cos ( 180 - q ) = -1* Cos q + 0 therefore Cos (180 - q) = -cos q

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

a. sin (a -b) = sin a cos b – cos a sin b b. cos (a -b) = cos a cos b – sin a sin b c. cos (a -b) = cos a cosb + sin a sin b d. sin (a + b) = sin a cos b + cos a sin b

OpenStudy (anonymous):

so d will be the answer

OpenStudy (butterflydreamer):

are you sure it's d? (hint: look at what i used... i used cos ( a - b) = cos a cos b + sin a sin b )

OpenStudy (anonymous):

ok i got it now thanks

OpenStudy (butterflydreamer):

you're welcome :)